​Betsy, a recent​ retiree, requires ​$6,000 per year in extra income. She has ​$50,000 to invest and can invest in​ B-rated bonds paying 13​% per year or in a certificate of deposit​ (CD) paying 3​% per year. How much money should be invested in each to realize exactly ​$6,000 in interest per​ year?
Answer by Theo(13342)

let x = the amount invested in b rated bonds.
let y = the amount invested in certificates of deposit.

x + y = 50000

this means that the total investment must be equal to 50,000 dollars.

.13x + .03y = 6000

this means that the total interest must be equal to 6,000 dollars.

these are 2 equations that need to be solved simultaneously.

i will solve by elimination.

start with:

x + y = 50000 (first equation)
.13x + .03y = 6000 (second eauation)

multiply both sides of the first equation by 13 and multiply both sides of the second equation by 100 to get:

13x + 13y = 650000
13x + 3y = 600000

subtract the second equation from the first to get 10y = 50000.

divide both sides of that equation by 10 to get y = 5000

since x + y = 50000, then x must be equal to 45000.

your solution should be that 45000 must be invested at 13% and 5000 must be invested at 3%.

.13 * 45000 + .03 * 5000 = 6000.

the solution looks good.